please help me solve the quadratic equation
12x^2 + 36a^2 = 43ax

Answer by ankor@dixie-net.com(22740)

12x^2 + 36a^2 = 43ax
:
Arrange in the general form:
12x^2 - 43ax + 36a^2 = 0
:
This will factor to:
(4x - 9a)(3x - 4a) = 0
:
4x = 9a
x = (9/4)a
and
3x = 4a
x = (4/3)a